OpenStudy (christos):

What is the dy/dx of 1/x ?

hartnn (hartnn):

use the standard formula for x^n 1/x = x^(-1)

OpenStudy (anonymous):

\[1/x= x^{-1}\] \[dy/dx x^{-1}=-x^{-2}\] Bring down the exponent and then subtract one from it :)

OpenStudy (christos):

So according to yyou this is wrong? http://screencast.com/t/C1ovmjAgq7b

hartnn (hartnn):

ofcourse its correct, you got the result -1/x^2, correct. just the method used is different

OpenStudy (anonymous):

Don't forget that a negative (-) exponent can be removed, and the term flipped. They said \[-1/x^{2}\]which equals my answer of \[-x^{−2}\] Raising something to a negative exponent basically just flips it :) And @hartnn is right as well, they did it with the limit definition of a derivative, rather than shortcuts

OpenStudy (christos):

I see!! Thanks a lot !!